OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 3) and is perpendicular to the line y = –x – 3?

OpenStudy (anonymous):

is it I assume you mean parallel and perpendicular to y = -2x though since y=-2 is a horizontal line passing through (0,-2) , (1,-2) and so on Parallel: y = -2x-3 Perpendicular y = 0.5x-3

OpenStudy (anonymous):

but, my options are: y = –x – 1 y = –x + 5 y = x + 1 y = x + 5

jimthompson5910 (jim_thompson5910):

what is the slope of the line y = –x – 3